\providecommand{\bysame}{\leavevmode\hbox to3em{\hrulefill}\thinspace}
\providecommand{\MR}{\relax\ifhmode\unskip\space\fi MR }
% \MRhref is called by the amsart/book/proc definition of \MR.
\providecommand{\MRhref}[2]{%
  \href{http://www.ams.org/mathscinet-getitem?mr=#1}{#2}
}
\providecommand{\href}[2]{#2}
\begin{thebibliography}{VDMZ00}

\bibitem[AFG24]{asselah2024critical}
Amine Asselah, Nicolas Forien, and Alexandre Gaudilli\`ere, \emph{The critical
  density for activated random walks is always less than 1}, Ann. Probab.
  \textbf{52} (2024), no.~5, 1607--1649. \MR{4791417}

\bibitem[BGH18]{basu2018nonfixation}
Riddhipratim Basu, Shirshendu Ganguly, and Christopher Hoffman,
  \emph{Non-fixation for conservative stochastic dynamics on the line}, Comm.
  Math. Phys. \textbf{358} (2018), no.~3, 1151--1185. \MR{3778354}

\bibitem[CGHL17]{candellero2017oil}
Elisabetta Candellero, Shirshendu Ganguly, Christopher Hoffman, and Lionel
  Levine, \emph{Oil and water: a two-type internal aggregation model}, Ann.
  Probab. \textbf{45} (2017), no.~6A, 4019--4070. \MR{3729622}

\bibitem[DRS10]{dickman2010activated}
Ronald Dickman, Leonardo~T. Rolla, and Vladas Sidoravicius, \emph{Activated
  random walkers: facts, conjectures and challenges}, J. Stat. Phys.
  \textbf{138} (2010), no.~1-3, 126--142. \MR{2594894}

\bibitem[DVZ98]{dickman1998self}
Ronald Dickman, Alessandro Vespignani, and Stefano Zapperi,
  \emph{Self-organized criticality as an absorbing-state phase transition},
  Phys. Rev. E \textbf{57} (1998), no.~5, 5095.

\bibitem[FMR09]{Fey2007StabilizabilityAP}
Anne Fey, Ronald Meester, and Frank Redig, \emph{Stabilizability and
  percolation in the infinite volume sandpile model}, Ann. Probab. \textbf{37}
  (2009), no.~2, 654--675. \MR{2510019}

\bibitem[For25]{forien2025newproof}
Nicolas Forien, \emph{A new proof of superadditivity and of the density
  conjecture for {A}ctivated {R}andom {W}alks on the line}, arXiv:2502.02579,
  2025.

\bibitem[HJJ24]{hoffman2024density}
Christopher Hoffman, Tobias Johnson, and Matthew Junge, \emph{The density
  conjecture for activated random walk}, arXiv:2406.01731, 2024.

\bibitem[HJJ25]{hoffman2025hockey}
\bysame, \emph{The hockey-stick conjecture for activated random walk},
  arXiv:2411.02541, 2025.

\bibitem[HJJM25]{hoffman2025cutoff}
Christopher Hoffman, Tobias Johnson, Matthew Junge, and Josh Meisel,
  \emph{Cutoff for activated random walk}, arXiv:2501.17938, 2025.

\bibitem[Hoe63]{hoeffding1963probability}
Wassily Hoeffding, \emph{Probability inequalities for sums of bounded random
  variables}, J. Amer. Statist. Assoc. \textbf{58} (1963), 13--30. \MR{144363}

\bibitem[HRR23]{hoffman2023active}
Christopher Hoffman, Jacob Richey, and Leonardo~T. Rolla, \emph{Active phase
  for activated random walk on {$\mathbb {Z}$}}, Comm. Math. Phys. \textbf{399}
  (2023), no.~2, 717--735. \MR{4576759}

\bibitem[Hu22]{hu2022active}
Yiping Hu, \emph{Active phase for activated random walk on $\mathbb{Z}^2$},
  arXiv:2203.14406, 2022.

\bibitem[JKM25]{junge2025meanfield}
Matthew Junge, Harley Kaufman, and Josh Meisel, \emph{Mean-field limit for
  activated random walk on the integer lattice}, arXiv:2509.10616, 2025.

\bibitem[LPS16]{levine2016apollonian}
Lionel Levine, Wesley Pegden, and Charles~K. Smart, \emph{Apollonian structure
  in the {A}belian sandpile}, Geom. Funct. Anal. \textbf{26} (2016), no.~1,
  306--336. \MR{3494492}

\bibitem[LPS17]{levine2017apollonian}
\bysame, \emph{The {A}pollonian structure of integer superharmonic matrices},
  Ann. of Math. (2) \textbf{186} (2017), no.~1, 1--67. \MR{3664999}

\bibitem[LS21]{levine2021howfar}
Lionel Levine and Vittoria Silvestri, \emph{How far do activated random walkers
  spread from a single source?}, J. Stat. Phys. \textbf{185} (2021), no.~3,
  Paper No. 18, 27. \MR{4334780}

\bibitem[LS24]{levine2024universality}
\bysame, \emph{Universality conjectures for activated random walk}, Probab.
  Surv. \textbf{21} (2024), 1--27. \MR{4718500}

\bibitem[PS13]{pegden2013convergence}
Wesley Pegden and Charles~K. Smart, \emph{Convergence of the {A}belian
  sandpile}, Duke Math. J. \textbf{162} (2013), no.~4, 627--642. \MR{3039676}

\bibitem[Rol20]{rolla2020activated}
Leonardo~T. Rolla, \emph{Activated random walks on {$\mathbb{Z}^d$}}, Probab.
  Surv. \textbf{17} (2020), 478--544. \MR{4152668}

\bibitem[RS12]{rolla2012absorbing}
Leonardo~T. Rolla and Vladas Sidoravicius, \emph{Absorbing-state phase
  transition for driven-dissipative stochastic dynamics on {${\mathbb{Z}}$}},
  Invent. Math. \textbf{188} (2012), no.~1, 127--150. \MR{2897694}

\bibitem[RSZ19]{rolla2019universality}
Leonardo~T. Rolla, Vladas Sidoravicius, and Olivier Zindy, \emph{Universality
  and sharpness in activated random walks}, Ann. Henri Poincar\'{e} \textbf{20}
  (2019), no.~6, 1823--1835. \MR{3956161}

\bibitem[ST17]{sidoravicius2017absorbing}
Vladas Sidoravicius and Augusto Teixeira, \emph{Absorbing-state transition for
  stochastic sandpiles and activated random walks}, Electron. J. Probab.
  \textbf{22} (2017), Paper No. 33, 35. \MR{3646059}

\bibitem[ST18]{stauffer2018critical}
Alexandre Stauffer and Lorenzo Taggi, \emph{Critical density of activated
  random walks on transitive graphs}, Ann. Probab. \textbf{46} (2018), no.~4,
  2190--2220. \MR{3813989}

\bibitem[VDMZ00]{vespignani2000absorbing}
Alessandro Vespignani, Ronald Dickman, Miguel~A. Mu{\~n}oz, and Stefano
  Zapperi, \emph{Absorbing-state phase transitions in fixed-energy sandpiles},
  Phys. Rev. E \textbf{62} (2000), no.~4, 4564.

\end{thebibliography}

\documentclass{amsart}

\usepackage{amssymb}
\usepackage{mathrsfs}
\usepackage[stretch=10,shrink=10]{microtype}
\usepackage[showframe=false, margin = 1.5 in]{geometry}
\usepackage{mathtools}
\usepackage[shortlabels]{enumitem}
\usepackage{ifthen}
\usepackage{stmaryrd}
\usepackage[hidelinks]{hyperref}

\newcommand{\E}{\mathbf{E}}

\renewcommand{\P}{\mathbf{P}}
\newcommand{\1}{\mathbf{1}}

\DeclarePairedDelimiter\floor{\lfloor}{\rfloor}%
\DeclarePairedDelimiter\ceil{\lceil}{\rceil}%
\DeclarePairedDelimiter\ii{\llbracket}{\rrbracket}%
\newcommand{\bigmid}{\;\big\vert\;}

\newcommand{\ZZ}{\mathbb{Z}}

\newcommand{\eosos}[1]{\mathcal{O}_{#1}}
\newcommand{\ip}[1]{\mathcal{I}_{#1}}

\newcommand{\Ber}{\mathrm{Bernoulli}}
\newcommand{\Bin}{\mathrm{Bin}}
\newcommand{\Ee}{\mathscr{E}}
\makeatletter

\newcommand{\crist}[1][\@nil]{%
\def\tmp{#1}%
   \ifx\tmp\@nnil
       \rho_*
    \else
       \rho_*^{(#1)}
    \fi}
\makeatother

\newcommand{\Left}{\ensuremath{\mathtt{left}}}
\newcommand{\Right}{\ensuremath{\mathtt{right}}}
\newcommand{\Sleep}{\ensuremath{\mathtt{sleep}}}
\newtheorem{thm}{Theorem}
\newtheorem{lemma}[thm]{Lemma}
\newtheorem{prop}[thm]{Proposition}
\newtheorem{cor}[thm]{Corollary}
\newtheorem{conjecture}[thm]{Conjecture}
\newtheorem*{conjecture*}{Density Conjecture}

\theoremstyle{remark}
\newtheorem{remark}[thm]{Remark}
\theoremstyle{definition}

\newcommand{\rt}[2]{\mathcal{R}_{#2}(#1)}
\newcommand{\lt}[2]{\mathcal{L}_{#2}(#1)}

\newcommand{\m}{\mathfrak{m}}

\begin{document}

\title{The odometer in subcritical activated random walk}
\author{Tobias Johnson}
\address{Department of Mathematics, College of Staten Island, City University of New York}
\email{\texttt{tobias.johnson@csi.cuny.edu}}
\author{Jacob Richey} 
\address{The Alfr\'ed R\'enyi Institute of Mathematics, Budapest, Hungary}
\email{\texttt{jfrichey001@gmail.com}}

\begin{abstract}
  We consider the activated random walk particle system, a model of self-organized criticality,
  on $\ZZ$ with i.i.d.-Bernoulli initial configuration. We show that at subcritical
  density, the system's odometer function, which counts
  the number of actions taken at each site, has a stretched exponential tail.
  It follows that the expected odometer at each site is finite.
\end{abstract}

\maketitle

\section{Introduction} 
	Activated random walk (ARW) is an interacting particle system which exhibits the trademark properties of self-organized criticality. Inspired by various toy models of self-organized criticality, ARW was introduced in \cite{dickman2010activated} as a stochastic variant of the well-known abelian sandpile model. ARW has proven to be the most robust and tractable of these various models of self-organized criticality: in a nutshell, it is robust because it is random, and it is tractable because of the abelian property. 
	
	ARW evolves in the following way (see Section~\ref{sec:setup} and \cite{rolla2020activated} for more details). Particles are placed on a graph and can be in one of two states: active or sleepy. An active particle performs simple random walk and falls asleep at a given rate $\lambda \in (0,\infty)$ when alone on a site;
  a sleepy particle gets woken up when an active particle jumps to its location. At each site $x$ on the graph, we record the number of times that something occurs at $x$ (either a particle jumping away from $x$ or a particle falling asleep at $x$) by a function $m(x)$ called the odometer.
  This paper is concerned with the \emph{fixed-energy} version of ARW, which takes place on the graph
  $\ZZ^d$ with a translation-ergodic initial configuration of active particles
  whose density $\rho$ is a parameter of the system.
  A basic fact about ARW is that there exists a critical density $\rho_c = \rho_c(\lambda)$ such that the odometer is almost surely finite everywhere if $\rho < \rho_c$ and almost surely infinite everywhere if $\rho > \rho_c$ \cite{rolla2012absorbing,rolla2019universality}.
  
  Recently there have been significant strides toward establishing that ARW exhibits
  the hallmark properties of self-organized criticality.
  While fixed-energy sandpile models have a traditional absorbing-state phase transition in the parameter $\rho$,
  the \emph{driven-dissipative} versions of sandpile models were believed to drive themselves
  to a critical state. These systems take place on a finite graph with particles annihilated
  at the boundary. Each time the system stabilizes, with all particles sleepy or annihilated,
  a new particle is injected into the system to reactivate it.
  Dickman, Mu\~noz, Vespignani, and Zapperi explained that the critical state
  emerges because the driven-dissipative dynamics naturally push the system
  to the critical density of their fixed-energy version \cite{dickman1998self,vespignani2000absorbing}.
  This so-called \emph{density conjecture} was recently proven in dimension one for ARW \cite{hoffman2024density,forien2025newproof}.
  
  In this paper, we apply the tools developed in \cite{hoffman2024density} to analyze subcritical fixed-energy
  ARW on $\ZZ$. As we mentioned, below criticality the system \emph{fixates}, i.e., the amount of activity
  at each site is finite over all time. Because the system is abelian,
  under the usual sitewise representation we may view the actions of the particles as being
  determined by stacks of instructions at the sites of the graph (see Section~\ref{sec:setup}).
  The odometer thus encodes the system's final state, making it a central object of study.
  While much effort
  has gone into establishing the existence of nontrivial sub- and supercritical phases for ARW
  and bounding the critical densities \cite{rolla2012absorbing,sidoravicius2017absorbing,stauffer2018critical,basu2018nonfixation,hoffman2023active,hu2022active,asselah2024critical,junge2025meanfield},
  nothing is known about the odometer beyond its almost-sure finiteness below criticality.
  
  To state our results, let $m$ be the odometer for ARW on $\ZZ$ 
  (to be defined rigorously in Section~\ref{sec:setup}), 
  and let $\P_\rho$ and $\E_\rho$ denote probabilities
  and expectations with respect to i.i.d.-$\Ber(\rho)$ initial configurations of active particles.
  Our first result gives the tail behavior of the odometer below criticality:
\begin{thm}\label{thm:tailbound}
  For any sleep rate $\lambda>0$ and 
  $\rho<\rho_c(\lambda)$, there exist constants $c,\,c',\,C,\,C'>0$ depending
  only on $\lambda$ and $\rho$ such that for all $n$,
  \begin{align*}
    Ce^{-c\sqrt{n}}\leq\P_\rho\bigl(m(0)\geq n\bigr) \leq C'e^{-c'\sqrt{n}}.
  \end{align*}
\end{thm}
Note that our result is stated for the odometer at site~$0$ but holds at all sites
by translation invariance.
The upper-tail bound immediately proves:
\begin{cor}\label{cor:finite.expectation}
  For any sleep rate $\lambda>0$ and density $\rho<\rho_c(\lambda)$, we have $\E_\rho\bigl[ m(0) \bigr]<\infty$.
\end{cor}
This result represents the first progress toward
the question posed by Rolla of determining the conditions that give
the odometer finite expectation \cite[p.~517]{rolla2020activated}.

	Previous works on the abelian sandpile model \cite{pegden2013convergence,levine2016apollonian,levine2017apollonian} and on a similar \emph{oil and water} model \cite{candellero2017oil} have given precise descriptions of the scaling limit of the odometer function, but in the single-source setting, where the initial condition is a large stack of particles at the origin. Roughly speaking, this version represents a supercritical system relaxing to a stable state. The fixed-energy version we consider, where particles are initially distributed according to i.i.d.\ measure in infinite volume, differs significantly in that the background particle density is always subcritical. The results nearest to ours in the existing ARW literature appear in \cite{levine2021howfar}, where the authors consider a weaker notion of critical density based on finiteness of the third moment of $m(0)$ and prove some connections between it and various other critical densities (many of which are now known to be equivalent). For the abelian sandpile model in $\mathbb{Z}^d$ at sufficiently small background density (and in $d = 1$ at any subcritical density), it follows from \cite[Theorem~4.2]{Fey2007StabilizabilityAP} that the odometer has subexponential tail.  
  
  Our results raise the question of what happens at criticality.
  It is generally believed that fixed-energy ARW at criticality
  remains active forever, making the odometer function infinite at every site
  a.s.\ (see \cite[Section~1.3]{rolla2020activated}).
  This is thought to be a difficult problem, but the following conjecture
  seems more approachable:
  \begin{conjecture}\label{conj:main} For any sleep rate $\lambda>0$,
    as $\rho \nearrow \rho_c(\lambda)$, 
    we have $\E_\rho\bigl[ m(0)\bigr] \nearrow \infty$.
  \end{conjecture}
  Our lower bound in Theorem~\ref{thm:tailbound} is not strong enough to prove this conjecture.
   We believe that new tools are needed to understand
  ARW near and at criticality (see Remark~\ref{rmk:obstacle}).
	
	Another question raised by our results is whether they hold with initial conditions other
  than the i.i.d.-$\Ber(\rho)$ considered here. Indeed, ARW is believed to possess some
  \emph{universality}, meaning that its behavior should not depend on minor details of the model,
  and that it should typify a broader class of processes exhibiting
  self-organized criticality (see \cite{levine2024universality}). One example of its universality is the density
  conjecture in dimension one, which demonstrates that two different versions of ARW have the same
  critical density; in fact two additional variants of the model are also shown
  to have the same critical density \cite{hoffman2024density}. Another is
  the result of Rolla, Sidoravicius, and Zindy that the critical density for fixed-energy ARW
  on a lattice is the same for all translation-ergodic initial distributions with the same
  density \cite{rolla2019universality}.
  While our results should hold with nearly verbatim proofs for other i.i.d.\ distributions
  with light tails, the tail behavior of the odometer
  cannot be universal: the initial number of particles at a site is a lower bound on the odometer
  at that site, and hence for any initial distribution with heavier tails than $e^{-c\sqrt{n}}$,
  the odometer will have a heavier tail than in Theorem~\ref{thm:tailbound}.
  In the other direction, one could consider initial configurations that are even more regular
  than i.i.d., such as periodic ones. In this setting the proof of the upper bound
  in Theorem~\ref{thm:tailbound} should still work, but the proof of the lower bound will not,
  since it uses the fluctuations of the i.i.d.\ configuration in a crucial way.
  
  As for the universality of Corollary~\ref{cor:finite.expectation}, the expected odometer
  will be infinite even at subcriticality if the initial configuration has infinite variance.
  To see this, first note that if $n$ particles on a single site are stabilized,
  then the odometer at that site will be on the order of $n^2$ with high probability;
  this can be proven by the method used in Proposition~\ref{prop:supercritical}.
  Hence if $m$ is the odometer and $\sigma$ is the initial configuration, we have
  $\E[ m(0)]\geq C\E[\sigma(0)^2]$. We speculate that this lower bound might be sharp,
  and we (tentatively) conjecture that the odometer has finite expectation 
  if and only if the number of particles
  per site has finite variance.

  \subsection{Notation}
  We say that an event $E_n$ holds \emph{with overwhelming probability} (w.o.p.)\ if
  $\P(E_n)\geq 1-Ce^{-cn}$ for some constants $c,C>0$ with no dependence on $n$.
  Typically we will index w.o.p.\ events by $n$, but sometimes we will replace $n$
  with some other indexing variable $j$ and say
  that an event holds w.o.p.-$j$ to indicate that the bound is of the form $1-Ce^{-cj}$.
  The constants in w.o.p.\ bounds may depend on other parameters, but any such dependence
  will be specified when the notation is used. 
  We write $\ii{a,b}$ for the interval of integers $[a,b]\cap\ZZ$.
  


\section{Setup}\label{sec:setup} 

As in nearly all work on ARW, we use the \emph{sitewise representation}.
We give our notation and a brief overview of the construction, and refer the reader to 
\cite[Section~2]{rolla2020activated} for the details.
Each site on the underlying graph is given a stack of instructions.
If an active particle is present at a site, then the site may legally \emph{topple},
executing the next instruction on the stack.
The instruction tells the particle either to jump to a given neighboring site,
or to sleep (which is carried out only if no other particle is present at the site).
The value of this representation is that all sequences of legal topplings proceeding until
all particles are stabilized produce the same final configuration.
To state this formally, define
the \emph{odometer} given by a sequence of topplings as the function on the vertices of the graph
counting the number of topplings that occur at each site.
According to the \emph{abelian property} \cite[Lemma~2.4]{rolla2020activated}, 
all legal sequences of topplings within $V$ that produce a stable configuration on $V$ have the same odometer,
which we call the \emph{stabilizing odometer} for a given initial configuration
on a finite set of vertices $V$.
By taking limits, one can also define the stabilizing odometer on an infinite
set of vertices \cite[eq.~(2.6)]{rolla2020activated}, and in our setting of translation-ergodic
initial configurations of particles on lattices, fixation
of ARW occurs if and only if the stabilizing odometer for the entire graph is finite a.s.\ at all
sites \cite[Theorem~2.7]{rolla2020activated}. We also note that the final configuration
after stabilizing can be read off the stabilizing odometer (given
the initial configuration and instruction stacks) by observing the net flow
between neighboring particles. Thus, the odometer stabilizing vertices $V$
satisfies a mass-balance equation at each site that places either zero or one particle
at each site. To state the mass-balance equation, we restrict ourselves to the one-dimensional setting,
denoting the three types of instructions by \Sleep, \Left, and \Right. 
For a nonnegative integer--valued function $u$ on $\ZZ$, define $\lt{u}{k}$
(resp.\ $\rt{u}{k}$) as the number
of \Left\ (resp.\ \Right) instructions within the first $u(k)$ instructions at site~$k$.
We think of $\lt{u}{k}$ and $\rt{u}{k}$ as the number of \Left\ and \Right\ instructions executed
by the odometer $u$ at site~$k$. If $u$ is the odometer stabilizing initial configuration
$\sigma$ on vertices $V$, then for all $v\in V$, the function $u$ satisfies the mass-balance equation
\begin{align}\label{eq:mass.balance}
  \sigma(v) + \rt{u}{v-1} + \lt{u}{v+1} - \lt{u}{v}-\rt{u}{v} = \1\{\text{the $u(v)$th instruction
  at site~$k$ is \Sleep}\}.
\end{align}
Put another way, the odometer places zero or one particles at each site, and it places one particle
exactly when the final instruction executed at that site is \Sleep.

Functions other than the stabilizing odometer that satisfy \eqref{eq:mass.balance} are important
as well. We refer to any function $u\colon\ZZ\to\ZZ_{\geq 0}$ as an \emph{odometer}.
If it takes the value zero outside vertices $V\subseteq\ZZ$, we call it an \emph{odometer on $V$}.
And if it satisfies \eqref{eq:mass.balance} for all $v\in V$, we say that it is \emph{stable on $V$
for initial configuration $\sigma$}.
(Note that these two sets need not be the same; we will often consider
odometers on $\ii{0,n}$ stable on $\ii{1,n-1}$.)
The \emph{least-action principle} \cite[Lemma~2.3]{hoffman2024density} states that the stabilizing odometer
on $V$ is minimal among all odometers stable on $V$. Thus, the typical strategy
to estimate the stabilizing odometer is to give upper bounds by producing
a stable odometer and applying the least-action principle, and to give
lower bounds by finding some obstruction to the existence of a stable odometer taking small values.

In dimension one, a process called \emph{layer percolation} developed
in \cite{hoffman2024density} gives us some tools
for showing the existence and nonexistence of stable odometers.
We will describe the basic picture here and state some terminology, but we refer
the reader to \cite{hoffman2024density} for an introduction to the theory.
Layer percolation is a sort of $(2+1)$-dimensional directed percolation.
We think of it as a collection of \emph{cells} in dimension two, each of which infects
other cells in the next time step. We refer to a cell by $(r,s)_k$ for nonnegative integers $r$, $s$,
$k$, viewing $r\geq 0$ as its \emph{row}, $s\geq 0$ as its \emph{column}, 
and $k$ as the time step.
Then we write $(r,s)_k\to(r',s')_{k+1}$ to denote that cell $(r,s)_k$ infects
cell $(r',s')_{k+1}$.

Layer percolation provides a different representation of the set of stable odometers.
ARW can be coupled with layer percolation so that odometers on
$\ii{0,n}$ stable on $\ii{1,n-1}$ correspond to length~$n$ infection 
paths in layer percolation, i.e., a sequence of infections
\begin{align*}
  (0,0)_0=(r_0,s_0)_0\to(r_1,s_1)_1\to\cdots\to (r_n,s_n)_n.
\end{align*}
Roughly speaking, $r_k$ indicates the number of right-jump instructions
executed by the corresponding odometer at site~$k$, while $s_k$ indicates the number
of particles that the odometer leaves sleeping on $\ii{1,k}$ according to its
mass-balance equations. We then study infection paths in place of odometers, taking
advantage of subadditivity and of connections between layer percolation and branching processes.

We state this correspondence formally now. In \cite{hoffman2024density}, the instruction
stacks are extended to be two-sided. \emph{Extended odometers} that can take negative values
are defined, and the notion of odometers being stable at a site is generalized to these
extended odometers. (The details are unimportant, but executing negative quantities of instructions
corresponds to executing instructions in reverse, i.e., with a \Left\ instruction executed
at site $k$ pulling a particle from site~$k+1$ to site~$k$. Odometers taking negative
values do not have a useful meaning from the perspective of ARW, but allowing them
regularizes the set of odometers and permits the correspondence with infection paths.) We then 
define $\eosos{n}(\sigma,u_0,f_0)$ as the set of odometers $u$ on $\ii{0,n}$ that
\begin{enumerate}[(i)]
  \item are stable on $\ii{1,n-1}$ for initial configuration $\sigma$; \label{i:stable}
  \item satisfy $u(0)=u_0$;
  \item and satisfy $\rt{u}{0} - \lt{u}{1} = f_0$, i.e., induce a net flow of $f_0$ particles
    from site~$0$ to site~$1$.
\end{enumerate}

The class of extended odometers $\eosos{n}(\sigma,u_0,f_0)$ has a pointwise minimal element $\m$
that we call the \emph{minimal odometer}. Despite its name, it is in general only an extended odometer
and typically takes negative values at some sites. To construct it, set $\m(0)=u_0$
and then take $\m(1)$ to be the minimal value making $\rt{\m}{0} - \lt{\m}{1} = f_0$.
Then choose $\m(2)$ as the minimal value so that $\m$ is stable at site~$1$, choose
$\m(3)$ as the minimal value so that $\m$ is stable at site~$2$, and so on.

We are nearly ready now to state the correspondence between ARW and layer percolation.
For given $\sigma$, $u_0$, and $f_0$ (and instruction stacks, which we have suppressed from the notation),
we define a coupled instance of layer percolation \cite[Definition~4.4]{hoffman2024density}, 
and we write $\ip{n}(\sigma,u_0,f_0)$ to denote
the set of length~$n$ infection paths starting from cell $(0,0)_0$ in this instance of layer percolation.
We then obtain the following correspondence:
\begin{prop}[{\cite[Proposition~4.6]{hoffman2024density}}]\label{prop:4.6}
  Let $\m$ be the minimal odometer of $\eosos{n}(\sigma,u_0,f_0)$.
  Then there is a surjective map from $\eosos{n}(\sigma,u_0,f_0)$
  to $\ip{n}(\sigma,u_0,f_0)$ defined by taking $u\in\eosos{n}(\sigma,u_0,f_0)$ to the infection
  path $(r_0,s_0)_0\to\cdots\to(r_n,s_n)_n$ given by
  \begin{align*}
    r_j &= \rt{u}{j} - \rt{\m}{j},\\
    s_j &= \sum_{i=1}^j\1\{\text{instruction $u(i)$ at site~$i$ is \Sleep}\}.
  \end{align*}
  If two extended odometers $u,u'\in\eosos{n}(\sigma,u_0,f_0)$ map to the same infection path,
  then for each $j\in\ii{0,n}$ either $u(j)=u'(j)$ or $u(j)$ and $u'(j)$ are two indices
  in a string of consecutive \Sleep\ instructions at site~$j$.
\end{prop}
Thus $\eosos{n}(\sigma,u_0,f_0)$ and $\ip{n}(\sigma,u_0,f_0)$ are in bijection, if consecutive strings
of \Sleep\ instructions are collapsed.

The odometer stabilizing a region maximizes the number of particles sleeping in the region
(see \cite[Lemma~2.5]{hoffman2024density} for this dual version of the least-action
principle). Thus we are most interested in odometers corresponding to infection paths
ending in the highest possible row. While it is difficult to show that there
is a limiting density of particles after stabilizing larger and larger intervals in ARW,
it is easy to show that
\begin{align*}
  \frac1n \max\Bigl\{s_n\colon (r_0,s_0)\to\cdots\to(r_n,s_n)\in \ip{n}(\sigma,u_0,f_0)\Bigr\}
\end{align*}
has a deterministic limit which we call $\crist=\crist(\lambda)$ 
\cite[Proposition~5.18]{hoffman2024density}. Note that this limit does not depend
on $\sigma$, $u_0$, or $f_0$. It is natural to expect that $\crist=\rho_c$,
i.e., the limiting rate of growth in the second coordinate of layer percolation
is equal to the critical density for fixed-energy ARW on $\ZZ$, and indeed
this is one of the main results of \cite{hoffman2024density}.
When working with layer percolation, we will typically refer to $\crist$ rather than
$\rho_c$ since we are generally thinking of it in its guise as the limiting rate of
growth of the maximal infection path.

\section{Proofs}

To prove the upper bound on the tail of the odometer $m$, we try to construct
a stable odometer taking value $n$ at site~$0$ using layer percolation.
If we succeed, then $m(0)\leq n$ by the least-action principle. Thus, if we can show
that the probability of failure for our construction is $O(e^{-c\sqrt{n}})$, we obtain
our upper bound on $\P(m(0)>n)$. For numerical convenience,
we will instead construct a stable odometer taking value $3(1+\lambda)n^2$ at site~$0$ with
a failure probability of $O(e^{-cn})$.

To carry out the construction, we consider a greedy infection path in layer percolation corresponding to an extended odometer taking value $3(1+\lambda)n^2$ at site~$0$.
Then we need to confirm that this extended odometer takes nonnegative values, 
which we can do using branching process estimates. This proof follows the basic approach of \cite[Proposition~3]{hoffman2025hockey}, with a different set of estimates for proving nonnegativity.

For the lower bound, we consider the large deviation event that the initial configuration has density
greater than $\crist$ on an interval of width~$n$ around site~$0$.
We then seek to show that conditional on this event,
the odometer stabilizing the interval is $\Omega(n^2)$ at site~$0$ with high probability.
Since this large deviation event has probability
at least $e^{-cn}$ for some $c>0$, it will follow that $\P(m(0)> c'n^2)\geq Ce^{-cn}$,
thus proving the lower bound in Theorem~\ref{thm:tailbound}.
To prove the statement about the odometer,
we condition on the event and then use layer percolation to consider the 
class of stable extended odometers taking some value $u_0$ at site~$0$ with $u_0\leq n^2$.
Then, we want to show that all such extended odometers
take a negative value somewhere. Following the approach developed in 
\cite[Proposition~4.1]{hoffman2025cutoff},
this statement can be translated into
a bound on the growth of infection paths in an instance of layer percolation,
which again follows from branching process estimates.

\subsection{Upper bound on the tail}

Because of the role played by the minimal odometer
$\m$ in the correspondence between odometers and infection paths, we need estimates for $\rt{\m}{j}$.
The following lemma gives two estimates. In \ref{i:min1}, which holds for $j\leq n$,
our choice to set $\m(0)=3(1+\lambda)n^2$ is the dominant factor in the size of $\rt{\m}{j}$.
In \ref{i:min2}, we give a bound without dependence on $n$ which holds for all large $j$.

\begin{remark}\label{rmk:projective.limit}
  In the following lemma and beyond,
  we consider $\eosos{N}(\sigma,u_0,f_0)$
  and $\ip{N}(\sigma,u_0,f_0)$ for an arbitrarily large value of $N$ that
  plays no further role in the results. It would be conceptually simpler to define
  $\eosos{\infty}(\sigma,u_0,f_0)$ as the set of odometers on $\llbracket 0,\infty)$
  stable on $\llbracket 1,\infty)$ 
  and to define $\ip{\infty}(\sigma,u_0,f_0)$ as the set
  of infinite infection paths in layer percolation. The theory of \cite{hoffman2024density}
  generalizes with no trouble to this infinite setting---the infinite versions of all objects
  are the projective limits of their finite versions---but 
  we have chosen to avoid the work of setting up this formalism since the only gain
  is a slight simplification of notation.
\end{remark}

	\begin{lemma}\label{lem:min.odom}
    Let $\m$ be the minimal odometer of $\eosos{N}(\sigma,u_0,f_0)$
    with $f_0\in\{0,1\}$ and $u_0\geq 3(1+\lambda)n^2$. Suppose $\sigma$ consists of
    i.i.d.-$\Ber(\rho)$ particles at each site.
    \begin{enumerate}[(a)]
      \item\label{i:min1}
    For $j\leq n\leq N$,
    \begin{align*}
      \P\bigl( \rt{\m}{j}\geq n^2-\rho j^2/2 \bigr) &\geq 1 - Ce^{-cn}
    \end{align*}
    for $c,C>0$ depending only on $\lambda$ and $\rho$.
    \item\label{i:min2} Let $\epsilon>0$. For $n\leq j \leq N$,
    \begin{align*}
      \P\bigl( \rt{\m}{j}\geq -(\rho+\epsilon) j^2/2 \bigr)&\geq 1 - Ce^{-cj}
    \end{align*}
    for $c,C>0$ depending only on $\lambda$, $\rho$, and $\epsilon$.    
    \end{enumerate}
  \end{lemma}
  \begin{proof}
    In both cases, the result is a straightforward consequence
    of \cite[Proposition~5.8]{hoffman2024density},  a concentration bound for $\rt{\m}{j}$ around 
    \begin{align}\label{eq:5.8.center}
      \frac{u_0}{2(1+\lambda)} - \sum_{i=1}^j \biggl( f_0 + \sum_{v=1}^i\sigma(v)\biggr),
    \end{align}
    which for $u_0=3(1+\lambda)n^2$ is approximately equal to $\frac32 n^2 - \rho j^2/2$.

    We allow the constants in w.o.p.\ bounds to depend on $\lambda$ and $\rho$ now.
    Let $Z_i=\sum_{v=1}^i\sigma(v)$.
    By \cite[Proposition~5.8]{hoffman2024density}, for all $j\leq n$
    \begin{align}\label{eq:lowj1}
      \Biggl\lvert\rt{\m}{j}-
         \biggl(\frac{u_0}{2(1+\lambda)} - \sum_{i=1}^j ( f_0 + Z_i)\biggr)
         \Biggr\rvert &\leq n^2/100 \text{ w.o.p.-$n$.}
    \end{align}
    Since $Z_i\sim\Bin(i,\rho)$, for any $i\leq n$ we have
    $Z_i\leq\rho i + n/100$ w.o.p.-$n$ by Hoeffding's inequality.
    Thus by a union bound it holds w.o.p.-$n$ that
    $Z_i\leq\rho i + n/100$ for all $i\leq n$,
    which together with $u_0\geq 3(1+\lambda)n^2$ and $f_0\leq 1$ proves that
    \begin{align}
      \frac{u_0}{2(1+\lambda)} - \sum_{i=1}^j ( f_0 + Z_i)
      &\geq 1.5 n^2 - j - \rho j(j+1)/2 - jn/100 \text{ w.o.p.-$n$}\nonumber\\
        &\geq 1.4n^2-\rho j^2/2 \text{ w.o.p.-$n$}.
      \label{eq:lowj2}
    \end{align}
    Now \eqref{eq:lowj1} and \eqref{eq:lowj2} prove \ref{i:min1}.

    The proof of \ref{i:min2} is very similar.
    Let the constants in w.o.p.\ expressions depend on $\epsilon$
    in addition to $\lambda$ and $\rho$.
    By \cite[Proposition~5.8]{hoffman2024density}, for $j\geq n$,
    \begin{align}\label{eq:highj1}
      \Biggl\lvert\rt{\m}{j}-
         \biggl(\frac{u_0}{2(1+\lambda)} - \sum_{i=1}^j ( f_0 + Z_i)\biggr)
         \Biggr\rvert &\leq \epsilon j^2/4 \text{ w.o.p.-$j$.}
    \end{align}
    Using the bound that $Z_i\leq \rho i + \epsilon j/100$ for $i\leq j$ w.o.p.-$j$,
    \begin{align}
      \frac{u_0}{2(1+\lambda)} - \sum_{i=1}^j ( f_0 + Z_i)
      &\geq  - j - \rho j(j+1)/2 - \epsilon j^2/100 \text{ w.o.p.-$j$}\nonumber\\
        &\geq -(\rho+\epsilon/2) j^2/2  \text{ w.o.p.-$j$}.
      \label{eq:highj2}
    \end{align}
    Now \eqref{eq:highj1} and \eqref{eq:highj2} prove \ref{i:min2}.    
  \end{proof}
  
  Generally speaking, it is most useful to construct odometers putting a high density
  of particles to sleep. Such odometers correspond to infection paths in layer percolation
  that increase in their second coordinate
   at rate close to $\crist$, which can be constructed as \emph{greedy infection paths}
  (see \cite[Section~5.4]{hoffman2024density}).
  The following lemma gives the rate of growth of their first coordinate.  
  \begin{lemma}\label{lem:greedy.path}
    Fix $\epsilon>0$. Let $(0,0)_0\to(r_1,s_1)_1\to\cdots$ be the $k$-greedy infection
    path. Fix $\epsilon>0$.
    For constants $c$, $C$, and $k_0$ depending only on $\lambda$ and $\epsilon$,
    it holds for all $k\geq k_0$ and $j\geq 1$ that
    \begin{align*}
      \P\bigl( r_j \geq (\crist - \epsilon)j^2/2 \bigr) \geq 1 - Ce^{-cj}.
    \end{align*}
  \end{lemma}
  \begin{proof}
    This result follows immediately from \cite[Proposition~5.16]{hoffman2024density}.
    In detail, by \cite[Proposition~5.18]{hoffman2024density}
    we can choose $k_0=k_0(\lambda,\epsilon)$ 
    large enough that $\crist[k] \geq \crist-\epsilon/2$ for $k\geq k_0$.
    Now apply \cite[Proposition~5.16]{hoffman2024density} with $n$ in \cite[eq.~(38)]{hoffman2024density}
    replaced by $j$ and $t$ replaced by $\epsilon \sqrt{j}/4$.
  \end{proof}

  \begin{proof}[Proof of Theorem~\ref{thm:tailbound}, upper bound]
    Let $u_0=\ceil{3(1+\lambda)n^2}$. If $\sigma(0)=1$, then take $f_0=1$; otherwise
    let $f_0=0$. Let $N\geq n$, and
    consider the set of extended odometers $\eosos{N}(\sigma,u_0,f_0)$.
    Let $g$ be the extended odometer in this class
    corresponding to the $k_0$-greedy infection path, where $k_0$ is the constant
    from Lemma~\ref{lem:greedy.path} for $\epsilon=(\crist-\rho)/3$. 
    We will show that $g$ is a genuine odometer w.o.p.-$n$, i.e., it takes nonnegative values everywhere on
    $\ii{0,N}$. Here the constants in w.o.p.\ bounds may depend on $\lambda$ and $\rho$.
    By Lemma~\ref{lem:min.odom}\ref{i:min1},
    it holds w.o.p.-$n$ that $\m(j)\geq 0$ for all $0\leq j\leq n$,
    since $n^2-\rho j^2/2\geq 0$ for $j\leq n$. Hence by the minimality of $\m$ we have
    \begin{align*}
      g(j) \geq \m(j)\geq 0 \text{ for all $0\leq j\leq n$ w.o.p.-$n$}
    \end{align*}
    To show that $g(j)$ is likely to be positive for $j> n$,
    we apply Lemma~\ref{lem:min.odom}\ref{i:min2} together with Lemma~\ref{lem:greedy.path}
    to show that for $j>n$,
    \begin{align*}
      \P\Bigl( r_j+\rt{\m}{j} < (\crist-\rho-2\epsilon)j^2/2\Bigr) \leq Ce^{-cj},
    \end{align*}
    where $r_j$ is the first coordinate of the $k_0$-greedy infection path at step~$j$.
    According to Proposition~\ref{prop:4.6}, we have $g(j)=r_j+\rt{\m}{j}$.
    We have chosen $\epsilon$ small enough that $\crist-\rho-2\epsilon>0$,
    and now a union bound in $j$ shows that
    \begin{align*}
      g(j) \geq 0 \text{ for all $n\leq j\leq N$ w.o.p.-$n$}
    \end{align*}
    Thus, we have shown that with overwhelming probability, there exists
    a (nonnegative) odometer $g$ on $\ii{0,N}$ that is stable on $\ii{1,N-1}$,
    satisfies $g(0)=u_0$, and induces a net flow of $\sigma(0)$ particles from site~$0$ to site~$1$.
    
    By the same reasoning, with overwhelming probability there exists
    an odometer $g'$ on $\ii{-N,0}$ that is stable on $\ii{-N+1,1}$, satisfies $g'(0)=u_0$,
    and induces a net flow of zero particles from $0$ to $-1$ (here we are taking
    $f_0=0$ in all cases).
    Pasting $g$ and $g'$ together produces an odometer that leaves no particles
    at site~$0$, since it induces a net flow of $\sigma(0)$ particles from site~$0$ to site~$1$
    and of zero particles from site~$0$ to site~$-1$. Reducing the value of this pasted odometer
    at site~$0$ to the first non-\Sleep\ instruction before instruction~$u_0$ 
    (which exists w.o.p.-$n$) yields
    an odometer on $\ii{-N,N}$ stable on $\ii{-N+1,N-1}$.
    
    Now, let $m^{(N)}$ denote the odometer stabilizing configuration $\sigma$ on $\ii{-N+1,N-1}$.
    By the least-action principle, we have shown that
    \begin{align*}
      \P \Bigl(m^{(N)}(0)\geq \ceil[\bigl]{3(1+\lambda)n^2}\Bigr) \leq Ce^{-cn}
    \end{align*}
    for some constants $c,C>0$ depending only on $\lambda$ and $\epsilon$
    (which depends on $\rho$).
    Since $m(0)$ is the increasing limit of $m^{(N)}(0)$ as $N\to\infty$
    \cite[eq.~(2.6)]{rolla2020activated}, we have
    \begin{align*}
      \P \Bigl(m(0)\geq \ceil[\big]{3(1+\lambda)n^2}\Bigr) \leq Ce^{-cn}.
    \end{align*}
    And finally by substituting for $n$ and altering the constants we obtain the upper bound
    in Theorem~\ref{thm:tailbound}.
  \end{proof}
  
  \begin{remark}
    In the previous proof, we construct an odometer on $\ii{-N,N}$ stable on
    $\ii{-N+1,N-1}$ taking value $3(1+\lambda)n^2$ at site~$0$ with
    probability $1-Ce^{-cn}$. Since $N$ plays no role in the failure probability, the construction
    can continue indefinitely to form a stable odometer on $\ZZ$, in line with
    Remark~\ref{rmk:projective.limit}. Buried in the proof is the quirk that
    this odometer would stabilize the ARW system leaving a density of particles
    strictly greater than the initial density.
    (This fact is never stated explicitly in this paper, but the way that the proof works
    is to choose a greedy infection path whose growth rate in the second coordinate
    is greater than the initial density. By Proposition~\ref{prop:4.6}, this growth
    rate is the density that the corresponding odometer leaves behind.)
    The explanation for this apparent paradox
    is that the odometer grows as the distance from the origin increases, which has the effect
    of funneling particles in the system toward the origin. The true odometer, of course,
    is translation invariant and does not alter the density of the system.
  \end{remark}

  \subsection{Lower bound on the tail}\label{sec:lower}
  
  First, we prove our claim that conditional on a high initial density on a length~$n$
  interval, the stabilizing odometer at the origin is $\Omega(n^2)$.
  \begin{prop}\label{prop:supercritical}
    Let $\sigma$ be chosen uniformly from all configurations
    of active particles on $I:=\ii[\big]{-\ceil{n/2}+1,\,\floor{n/2}}$ containing exactly
    $\floor{(\crist+\epsilon)n}$ particles.
    Let $g$ be the stabilizing odometer on $I$ for initial configuration $\sigma$.
    For constants $c,C>0$ depending only on $\epsilon$ and $\lambda$,
    \begin{align*}
      \P\biggl(g(0) > \frac{(1+\lambda)\epsilon}{32} n^2\biggr) \geq 1 - Ce^{-cn}.
    \end{align*}
  \end{prop}
  \begin{proof}\newcommand{\altn}{\tilde{n}}
    Let $\altn=\ceil{n/2}-1$.
    The idea is to consider odometers on the interval $\ii{0,\altn}$, which will stand
    in for either the left or right half of $I$.
    Using layer percolation, we will argue that with high probability,
    all stable extended odometers on $\ii{0,\altn}$ that are
    below $(1+\lambda)\epsilon n^2/32$ at site~$0$ are negative
    at site~$\altn$. Hence all stable odometers, including the actual stabilizing odometer,
    are at least $(1+\lambda)\epsilon n^2/32$ at site~$0$.

    Let $\sigma'$ denote a configuration on $\ii{0,\altn}$ obtained by restricting
    $\sigma$ to a given subinterval of length~$\altn+1$ and then relabeling its sites as $0,\ldots,\altn$.
    Suppose that $0\leq f_0\leq n$ and $u_0 \leq 2\beta(1+\lambda)n^2$ for $\beta=\epsilon/64$.
    Our goal is to show that all extended odometers in $\eosos{\altn}(\sigma',u_0,f_0)$ take a negative value.
    We allow the constants in w.o.p.\ bounds to depend on $\lambda$ and $\epsilon$.
    
    Let $\m$ be the minimal odometer in $\eosos{\altn}(\sigma',u_0,f_0)$.
    We claim that
    \begin{align}
      \rt{\m}{\altn} \leq \biggl(\beta-\frac{\crist+\epsilon}{8} + \frac{\epsilon}{32}\biggr)n^2 
         \quad\text{w.o.p.}\label{eq:minnn.bound}
    \end{align}
    To prove this claim, let $Z_j=\sum_{i=1}^j\sigma'(i)$, and note that $\rt{\m}{\altn}$ is within
    $\epsilon n^2/64$ of 
    \begin{align*}
      \frac{u_0}{2(1+\lambda)} - \sum_{i=1}^{\altn}(f_0 + Z_i)
    \end{align*}
    with overwhelming probability by \cite[Proposition~5.8]{hoffman2024density}.
    Since $u_0 \leq 2\beta(1+\lambda)n^2$ and $f_0\geq 0$, 
    we can thus prove \eqref{eq:minnn.bound} by showing that
    \begin{align}
      \beta n^2 - \sum_{i=1}^{\altn}Z_i \leq
      \biggl(\beta-\frac{\crist+\epsilon}{8} + \frac{\epsilon}{64}\biggr)n^2 
         \quad\text{w.o.p.}\label{eq:migration.estimate}
    \end{align}
    The random variable $Z_i$ is the number of the $\floor{(\crist+\epsilon)n}$ particles
    placed randomly in an interval of length~$n$ that fall in a subinterval of length~$i$;
    that is, it is hypergeometrically distributed. Hoeffding's inequality
    gives the same tail bounds as for the distribution
    $\Bin\bigl(\floor{(\crist+\epsilon)n},\,i/n\bigr)$ \cite[Section~6]{hoeffding1963probability}.
    Hence $Z_i\geq (\crist+\epsilon)i+\epsilon n/64$ w.o.p. Taking a union bound over
    $i=1,\ldots,\altn$
    yields \eqref{eq:migration.estimate}, establishing \eqref{eq:minnn.bound}.
    
    By Proposition~\ref{prop:4.6}, for an extended odometer $g\in\eosos{\altn}(\sigma',u_0,f_0)$,
    the value of $\rt{g}{\altn}$ is given by $r_{\altn}+\rt{\m}{\altn}$,
    where $r_{\altn}$ is the column at step~$\altn$ of the corresponding infection
    path.
    Let
    \begin{align*}
      R_{\altn} = 
        \max\bigl\{r_{\altn}\colon (r_0,s_0)\to\cdots\to (r_{\altn},s_{\altn})\in\ip{\altn}(\sigma',u_0,f_0)\bigr\},
    \end{align*}
    the rightmost column infected starting from $(0,0)_0$ at step~$\altn$ of layer percolation.
    By \cite[Lemma~4.6]{hoffman2025cutoff},  we have $R_{\altn}\leq(\crist+\epsilon/2)n^2/8$ w.o.p.
    Using \eqref{eq:minnn.bound} to bound $\rt{\m}{\altn}$, 
    \begin{align*}
      R_{\altn}+\rt{\m}{\altn} &\leq \frac{\crist+\epsilon/2}{8}n^2 + 
      \biggl(\beta-\frac{\crist+\epsilon}{8} + \frac{\epsilon}{64}\biggr)n^2\quad\text{w.o.p}\\
      &\leq \bigl(\epsilon/16 + \beta - \epsilon/8 + \epsilon/64\bigr)n^2
        \leq -\epsilon n^2/32\quad\text{w.o.p.}
    \end{align*}
    Hence all extended odometers in $\eosos{\altn}(\sigma',u_0,f_0)$ are negative
    at site~$\altn$ w.o.p.
    
    To complete the proof, we now consider the left and right halves
    of the interval $I$. 
    Let $I_{-}:=\ii[\big]{-\ceil{n/2}+1,0}$, and let
    $I_+$ be defined as $\ii[\big]{0,\floor{n/2}}$ if $n$ is odd 
    or $\ii[\big]{0,n/2-1}$ if $n$ is even, so that $I_-$ and $I_+$
    are intervals of length~$\altn+1$.
    Applying our result on $\eosos{\altn}(\sigma',u_0,f_0)$ together
    with a union bound over the $O(\epsilon\lambda n^3)$
    choices of $u_0\leq 2\beta(1+\lambda)n^2$
    and $0\leq f_0\leq n$, we conclude that with overwhelming probability,
    if an odometer $g$ on $I_+$ is stable on $I_+\cap[1,\infty)$, then either
    $g(0) > 2\beta(1+\lambda)n^2$ or $g$ induces a strictly negative
    particle flow from site~$0$ to site~$1$; this stems from our (necessary) assumption
    that $f_0\geq 0$ for our result on $\eosos{\altn}(\sigma',u_0,f_0)$.
    By symmetry, it holds with overwhelming probability that
    if an odometer $g$ on $I_-$ is stable on $I_-\cap(-\infty,-1]$, then either
    $g(0) > 2\beta(1+\lambda)n^2$ or $g$ induces a strictly negative
    particle flow from site~$0$ to site~$-1$.
    
    Now, suppose $g$ is the odometer stabilizing $\sigma$ on $I$. 
    Since it is stable on
    $I_+\cap[1,\infty)$ and $I_-\cap(-\infty,-1]$, it holds with overwhelming
    probability that either $g(0)>2\beta(1+\lambda)n^2$ or
    $g$ induces strictly positive net flows both
    from site~$-1$ to site~$0$ and from site~$1$ to site~$0$.
    But this second alternative is impossible, since then $g$ would leave
    multiple particles at site~$0$ and fail to be stable there.
  \end{proof}

  \begin{proof}[Proof of Theorem~\ref{thm:tailbound}, lower bound]
    Choose some $\epsilon>0$ so that $\crist<\rho+\epsilon<1$, which we can do since
    $\crist<1$ for any sleep rate $\lambda>0$ by \cite{hoffman2023active}.
    Let $I=\ii[\big]{-\ceil{n/2}+1,\,\floor{n/2}}$, an interval of $n$ integers.
    Let $\sigma$ be the initial i.i.d.-$\Ber(\rho)$ configuration of particles on $\ZZ$,
    and define $\Ee_n$ as the event that $\sigma$ contains at least $\floor{(\rho+\epsilon)n}$ particles
    on $I$.
    
Take $\sigma'$ to be an empty configuration if $\Ee_n$ fails and otherwise to consist
    of exactly $\floor{(\rho+\epsilon)n}$ particles chosen uniformly at random from
    the particles of $\sigma$ in $I$. By exchangeability, given $\Ee_n$ the configuration
    $\sigma'$ is uniform over all collections of $\floor{(\rho+\epsilon)n}$ particles
    on $I$. By Proposition~\ref{prop:supercritical}, the odometer stabilizing  $\sigma'$
    given $\Ee_n$ is greater than $\alpha n^2$ at $0$ with probability $1 - Ce^{-cn}$,
    for $\alpha=(1+\lambda)(\rho+\epsilon-\crist)/32$.
    Since the stabilizing odometer on $I$ is increasing in the initial configuration,
    and the stabilizing odometer on $\ZZ$ is bounded from below by the stabilizing
    odometer on $I$, we have proven that
    \begin{align*}
      \P\bigl(m(0) > \alpha n^2\bigmid \Ee_n\bigr) \geq 1 - Ce^{-cn}
    \end{align*}
    for constants $\alpha$, $c$, and $C$ depending on $\rho$ and $\lambda$.
    In particular,
    \begin{align}\label{eq:odometer.big}
      \P(m(0)>\alpha n^2\mid\Ee_n) \geq 1/2,\qquad
    \end{align}
    for all $n\geq n_0(\rho,\lambda)$.
    
    By basic large deviations theory, we have $\P(\Ee_n)\geq Ce^{-cn}$ for some constants $c,C>0$
    depending on $\rho$ and $\crist$. Hence,
    \begin{align*}
      \P\bigl(m(0)>\alpha n^2\bigr)\geq Ce^{-cn}
    \end{align*}
    for some constants $\alpha,c,C>0$ depending on $\rho$ and $\lambda$.
    And this establishes the lower bound of Theorem~\ref{thm:tailbound}.
  \end{proof}
  
  \begin{remark}\label{rmk:obstacle}
    Using a more careful bound on the probability of event $\Ee_n$,
    one could prove that for constants $c=c(\lambda)$, $C=C(\lambda)$, and
    $n_0=n_0(\lambda,\rho)$, it holds for all $n\geq n_0$ that
    \begin{align*}
      \P\bigl( m(0) \geq n \bigr) &\geq Ce^{-c(\crist-\rho)^2\sqrt{n}}.
    \end{align*}
    In this version of the result, one can fix the sleep rate $\lambda$
    and then see that the asymptotic rate function
    \begin{align*}
      \limsup_{n\to\infty}-\frac{1}{\sqrt{n}}\log\P(m(0)\geq n)
    \end{align*}
    converges to $0$ as $\rho\nearrow\crist$. This result suggests that
    $\E m(n)\to\infty$ as $\rho\nearrow\crist$ but cannot prove it
    without some control on $n_0$.
    
    The reason we cannot control $n_0$
    is that the layer percolation machinery establishes supercritical behavior
    when stabilizing $(\crist+\epsilon)n$ particles on an interval of length~$n$
    for sufficiently large $n$ but gives no information on how large $n$ needs to be.
    It will require new ideas to remedy this deficiency.
    The ultimate source of it is that the proof \cite[Proposition~5.18]{hoffman2024density}
    that the layer percolation growth rate converges to $\crist$ is entirely qualitative,
    which stems from its use of subadditivity.
  \end{remark}

  \section*{Acknowledgments}
  T.J.\ was partially supported by NSF grant DMS-2503779 and gratefully acknowledges support from
Uppsala University and the Wenner--Gren Foundation. J.R.\ was partially supported by a Simons Foundation Targeted Grant 
awarded to the R\'enyi Institute. We thank the Erd\H{o}s Center for its hospitality
in hosting workshops where some of this research was carried out. We thank Ahmed Bou-Rabee
for pointing us to results giving bounds on the odometer for the abelian sandpile.
\bibliographystyle{amsalpha}
\bibliography{expectation}

\end{document}